A Bibliography of Publications about the MuPAD Symbolic Algebra Language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.11

Title word cross-reference

2W [Mur00]. D [Tsa00].

-ary [Mur00]. -ideal [Tsa00].

’01 [ACM01].

1.0 [GPS97]. 1.2.2 [Ano95b, Zim95]. 1.3 [Pos97]. 1.4 [Ano01g, Sor99]. 1994 [Hec94]. 1999 [WRL99]. 1st [Gro95].

2.0 [Ano01g, Ano01r]. 2003 [ACM03b].

5th [Car96].

’96 [Lak96]. ’97 [KH97, Tec98]. abstract [Ano01a, Thi08]. administration [Sor96]. adt [Ano01a]. Advanced [Har00b]. aided [Thi00a]. Algebra [Ano93a, Ano93b, Ano96, FGHK95, Got94, GPS97, uRH95, uRH96, MK94, OS96, WRL99, Wiw92, Wiw93, Ano01l, Ben98, Ces97, F^{+}93, Fuc94, GKW03, Klu96, LW99, LW00, Sor96, Gro95, Klu96, Car96]. Algebraic [ACM99, ACM03a, Glo98, GPS97, Gut04, Lak96, RLRML04, Thi00a, Thi00b, Thi08, Tra00, BK04, Ke00, KFC^{+}01]. Algebrasystem [Wiw93]. algorithm [Mur00, Pop00, Tsa96]. algorithms [Ano01s]. Amsterdam [Hec94]. analysis [Ke00]. Andrews [Tra00]. Annual [ACM01, ACM03b]. Anwendung [Ben98]. Anwendungen [Ano93b]. Apfel [SRRM^{+}97]. application [PDP00]. Applications [Gro95, WRL99, GKW03]. appraisement [BCK99]. Approach [ZP06]. approximation [Pop00]. Architectures [Ano93b]. Architekturen [Ano93b]. ary [Mur00]. August [ACM03a, FGHK95, Glo98, Tra00]. Austria [M^{+}95]. Ax [Ano01-35]. AXIOM [Ben98, Ben98]. Axioms [Ano01b, Ano01-35].

based [Nau96, Thi00b]. Basen [Got94]. Bases [GB97, RLRML04, Ano01i, DZ95, Got94, Sat00]. Basic [Ano01d]. BC [ACM99]. Behandlung [Ano93b]. Benutzerhandbuch [F^{+}93]. Bernoulli [Har00a]. Bielefeld [FGHK95]. Bildung [uFSGO95]. bis [SRRM^{+}97]. bivariate [KN00]. black [DK98]. Book [SRRM^{+}97]. Boolean [Sat00]. box [DK98]. Bührke [SRRM^{+}97].

California [ACM03b]. CALYPSO [Ces97]. Canada [ACM99]. Cantabria [Gut04]. Cat [Ano01-36]. Categories [Ano01b, Ano01-36]. Challenge [Tec98]. challenges [FM97]. checking [Ano01-40]. closure [Tsa00]. code [Sor96]. column [Vol97]. Combinat [Thi08, Ano01c]. Combinatorics [Thi08, Ano01c]. Commutative [GPS97]. Comparing [Fat03]. Comparison [LW99, LW00]. complete [Tsa96]. Complexes [Ano93b]. Computation [ACM03a, Bee17, KH97, Lak96, Tra00, BCK99, Ces97, Har00a, Hec94, KN00, Ke00, KFC^{+}01, Sat00, ACM99, Glo98, Gut04]. Computational [ACM01, ACM03b]. computations [AH09]. Computer [Ano93a, Ano93b, Got94, GKW03, GPS97, Gro95, uRH95, uRH96, MK94, OS96, WRL99, Wiw93, Ben98, Car96, Ces97, Klu96, LW99, LW00, Sor96, Thi00a, Thi00b, FGHK95, Klu96]. Computer-Algebra [Got94, Wiw93, Klu96]. Computer-Algebrasystem [Wiw93]. Computeralgebra [Ano93b, Ben98, uRH95, uRH96, Sor96]. Computeralgebra-Anwendungen [Ano93b]. Computeralgebra-Systemen [Ben98, Sor96]. Computing [Hub00, Maj02, Maj04, SW99]. Concept [Sor97b]. Conference [Gro95, M^{+}95, WRL99]. constraint [BK04]. Copernicus [Ano05].

Data [Wiw92, Ano01a, Ano01d, Ano01j, Ano01-29, F^{+}93, Fuc94, Ano96]. datatypes [Ano01d]. Demonstration [BG97]. denotational [Nau96]. denotationales [Nau96]. DERIVE [Ben98, Ben98]. detools [Ano01e]. development [Thi08, Klu96]. Developments [MO95]. Diego [ACM03b]. Differential [Ano93b, Ano01e, Tsa96, Got94]. Differential-Gröbner-Basen [Got94]. Differentialgleichungen [Ano93b]. Dimensional [ZP06]. distributed [Ber97, Nau97]. dla [Maj05b]. Dom [Ano01-37]. Domains [Ano01b, Ano01-37]. Dynamic [Sor97b, Sor97a, Sor97d, Sor97c, Sor99, Ano01q, Sor96]. Dynamische [Sor97c, Sor96].

ED [M^{+}95]. ED-MEDIA [M^{+}95]. Edition [mat91]. Education [uFSGO95, OS96, WKZZ03]. Educational [M^{+}95, Ano01-33]. efficiency [Sor96]. efficient [KFC^{+}01, LPP04]. Effizienz [Sor96]. Engineering [FGHK95, Ben98]. Entwicklung [Klu96]. enumeration [Tsa96]. environment [Klu96, SW99]. Equations [Ano93b, Ano01e]. error [Pop00]. essentials [Maj02, Maj04]. evaluation [KFC^{+}01]. Exact [RLRML04, LPP04]. Exchanging [BG97]. experiment [BK04]. Experimental [Ano93b]. Experimentelle [Ano93b]. exploration [Thi00b]. Expressions [BG97]. extended [Mur00].

factorizations [Tsa96]. fans [Hub00]. fast [Sor97c]. feedback [Thi08]. Flexibilität [Sor96]. flexibility [Sor96]. formatted [Ano01w]. foundations [GKW03]. Fourth [Ano05]. FOXBOX [DK98]. fp [Ano01f]. Fraser [ACM99]. Function [Bee17, Pop00]. functional [Ano01f]. future [DZ95].

Galilei [SRRM^{+}97]. Gambini [SRRM^{+}97]. Gauge [SRRM^{+}97]. GCD [Mur00]. generate [Ano01h, Ano01h]. geometric [KFC^{+}01]. Geometry [ACM01, ACM03b, GPS97, RLRML04]. German [Ano95c, F^{+}93, Got94, uRH95, uRH96, Klu96, Kra96, Nau96, Nau97, uFSGO95, Sor96, Sor97c, Wiw93, Ano93b, Ben98, SRRM^{+}97]. Germany [Glo98, FGHK95]. Geschlechterverhältnisse [SRRM^{+}97]. Getting [Maj05a]. graph [Ano01-38]. graphics [Ano01x]. graphs [Thi00a, Thi00b]. Gravity [SRRM^{+}97]. Gravity/MuPAD [SRRM^{+}97]. Graz [M^{+}95]. Grenzen [Sor97c]. Gröbner [Got94, Ano01i, DZ95, Got94, Hub00, RLRML04, Sat00]. groebner [Ano01i]. Group [mat91]. guide [Sor99].

Handbook [Bee17, F^{+}93, GKW03]. Haug [SRRM^{+}97]. HI [KH97]. High [SRRM^{+}97, SW99]. high-performance [SW99]. Higher [ZP06]. Higher-Dimensional [ZP06]. Human [Hec94]. Hybrid [KN00]. Hypermedia [M^{+}95].

ideal [Tsa00]. ideals [Hub00]. II [uRH96]. im [Got94]. IMACS [Gro95, WRL99]. impatient [Maj05b]. implementation [LPP04]. Implementing [Thi08]. import [Ano01j, Ano01j]. increase [Sor96]. independent [Oev00]. Ingenieurmathematik [Ben98]. input [Ano01h]. integer [Mur00]. integral [Ano01-34]. Integration [Sor97b, Sor97a, Sor97d, Ano01k, Kra96, Kra96]. interaction [Hec94]. International [ACM99, ACM03a, Glo98, Gro95, Gut04, KH97, Lak96, Tra00]. interpolation [KN00]. Intersecting [LPP04]. intersection [DLLP03]. intlib [Ano01k]. introduction [Oev00]. invariants [Thi00a, Thi00b]. Inverse [SRRM^{+}97, Pop00]. Involutive [GB97]. ISAAC [Gut04]. ISSAC [ACM99, ACM03a, Glo98, Lak96, Tec98, Tra00]. issues [Pos97]. iterative [Pop00].

Jauho [SRRM^{+}97]. Java [BCK99]. July [ACM99, Ano05, Gut04, KH97, Lak96]. June [ACM01, ACM03b, M^{+}95].

Kinetics [SRRM^{+}97]. Knots [SRRM^{+}97].

Level [ZP06]. Library [Bee17, Ano01a, Ano01c, Ano01e, Ano01f, Ano01i, Ano01n, Ano01m, Ano01l, Ano01q, Ano01o, Ano01-38, Ano01t, Ano01s, Ano01w, Ano01v, Ano01x, Ano01-28, Ano01z, Ano01-30, Ano01-32, Ano01-31, Ano01-34, Ano01-40, AH09, Ces97]. limits [Sor97c]. linalg [Ano01l]. linear [Ano01m, Ano01l, BK04, CCT04, Tsa96]. linopt [Ano01m]. Linux [uRH95, uRH96, uRH95, uRH96]. Liouvillian [CCT04]. Lise [SRRM^{+}97]. list [Ano01n]. listlib [Ano01n]. Loops [SRRM^{+}97]. Lübbig [SRRM^{+}97].

machine [Ber97, Sor96]. machine-code [Sor96]. MACSYMA [Ben98, Ben98, Tec98, PZ96b, PZ96a]. Magazine [mat91]. MAMMUT [Nau97]. management [Nau97]. manipulating [DK98]. manipulation [Ano01n, Ano01z, Ano01-32, Nau97, Nau97]. Manual [The96, Ano96, Sor99]. MAPLE [Ben98, Ben98, Ber97, PZ96b, PZ96a]. March [Hec94]. Maschinencode [Sor96]. Maschinencode-Objekte [Sor96]. Massachusetts [ACM01]. massively [Ber97]. matching [Ano01o]. matchlib [Ano01o]. Math [AH09, Ano95c]. MATHCAD [Ben98, Ben98]. MathCW [Bee17]. MATHEMATICA [Ben98, Ben98, Har00b, PZ96b, PZ96a, PDP00]. Mathematical [BG97, Bee17, Ano01-40]. Mathematical-Function [Bee17]. mathematics [Ben98, WKZZ03, uRH95, uRH96]. Mathematik [uRH95, uRH96]. Matheprobleme [Ano95c]. MATHLAP [Ben98, Ben98]. MathPAD [Ano05, mat91]. Matlab [AH09]. Maui [KH97]. Medford [ACM01]. MEDIA [M^{+}95]. Medizin [SRRM^{+}97]. Meeting [Ano05]. Meitner [SRRM^{+}97]. Meitner/Memel [SRRM^{+}97]. Memel [SRRM^{+}97]. memory [Ber97]. Minimal [GB97]. misc [Ano01p]. miscellanea [Ano01p]. model [Nau96]. Modell [Nau96]. Module [Sor96, Sor97c, Ano01q]. Modules [Sor97b, Sor97a, Sor97d, Ano01q, Sor96, Sor97c, Sor99]. MP [BG97]. multi [Fuc94, Ano96]. Multi-Processing [Ano96]. multidimensional [AH09].

Multimedia [M^{+}95]. multiplication [Fat03]. multiprecision [KFC^{+}01]. Multiprocessing [Wiw92, F^{+}93]. MuPAD [Ano93b, Ano05, Ben98, mat91, uRH95, uRH96, Maj05a, Ano93b, Ano95b, Ano95a, Ano01d, Ano01g, Ano01r, Ben98, CO04, DZ95, F^{+}93, Got94, HKM^{+}97, uRH95, uRH96, Maj02, Maj04, Maj05b, MK94, MO95, The96, Nau95, uFSGO95, Oev00, PZ96b, PZ96a, Pos97, SRRM^{+}97, Sor97b, Sor97a, Sor97d, Sor97c, Sor99, SW99, Thi08, Wiw92, Wiw93, Zim95, Ano96, Fuc94, McA99]. MuPAD-Tutorium [Oev00].

Naturwissenschaft [SRRM^{+}97]. Near [DLLP03]. Near-optimal [DLLP03]. nearly [Sor97c]. Network [Ano01-38]. Newtons [SRRM^{+}97]. Nicholas [Ano05]. nichtlinearen [Ano93b]. niecierpliwych [Maj05b]. nineteenth [ACM03b]. Non [CCT04]. Non-Liouvillian [CCT04]. Nonlinear [Ano93b]. notations [Har00b]. number [Ano01t]. numbers [Har00a]. numeric [Ano01s]. numerical [Ano01s]. numlib [Ano01t].

object [Nau96]. object-based [Nau96]. objects [DK98, Sor96]. objektbasierte [Nau96]. Objekte [Sor96]. ODE [PZ96b, PZ96a, Ano01u]. ODEs [CCT04]. ohne [Sor97c]. Online [mat91]. only [Thi08]. OpenMath [PDP00]. operator [Tsa96]. Optics [SRRM^{+}97]. Optimal [Pop00, DLLP03]. optimization [Ano01m]. order [CCT04]. ordinary [Tsa96]. oriented [LW99, LW00]. orthogonal [Ano01v]. orthpoly [Ano01v]. other [Ano01h]. output [Ano01w, Ano01w].

p [Ano93b]. PA [ACM03a]. Parallel [Ano93b, KH97, Sat00, Ber97, Ces97, Klu96, Nau96]. parallele [Ano93b, Nau96]. Parallelism [HKM^{+}97, Nau95]. Parallelverarbeitung [Klu96]. parameterization [DLLP03]. parametric [BK04]. PASCO [KH97]. pattern [Ano01o]. performance [SW99]. Perspectives [Ano93a]. Philadelphia [ACM03a]. Physik [SRRM^{+}97]. platform [Oev00]. play [Sor97c]. plot [Ano01x]. plotlib [Ano01y]. Poland [Ano05]. Polish [Maj05b]. Polmat [AH09]. polylib [Ano01z]. polynomial [Ano01z, Fat03, LW99, LW00]. polynomial-oriented [LW99, LW00]. polynomials [Ano01v]. Portable [Bee17]. PRECISE [KFC^{+}01]. predefined [Ano01-35, Ano01-36, Ano01-37]. predicates [KFC^{+}01]. Pref [Ano01-39]. preferences [Ano01-39]. Pro [Maj02, Maj04]. problem [Mur00, SW99, SRRM^{+}97]. Problem/Bührke [SRRM^{+}97]. Problems [FM97, Ano95c, Tec98]. Proceedings [ACM99, ACM03a, Glo98, Gut04, KH97, Tra00, Lak96, M^{+}95, ACM01, ACM03b, Car96]. Process [SRRM^{+}97]. Process/Haug [SRRM^{+}97]. processing [Fuc94, Klu96, Ano96]. Products [Lin97]. prog [Ano01-27]. Program [Ano93b, Ano95c]. Programm [Ano95c]. programmer [Ano01-27]. Programmierumgebung [Klu96]. Programming [Bee17, ZP06, Ano01f, BK04, Klu96, Sor99]. Programmpakete [Ano93b]. programs [Ano01h, Fat03]. Propagation [ZP06]. properties [Ano01-28, Ano01-40]. property [Ano01-28]. Pullin [SRRM^{+}97]. puzzles [FM97].

quadrics [DLLP03, LPP04]. Quantum [SRRM^{+}97, SRRM^{+}97]. Questioning [SRRM^{+}97]. Quick [Ano01r].

Rapid [Har00a]. rational [KN00]. Reference [Ano01r]. related [Har00a]. reliable [KFC^{+}01]. Renneberg [SRRM^{+}97]. report [Vol97]. representation [DK98]. Research [mat91]. Review [Ber96, PZ96b, PZ96a]. reviews [SRRM^{+}97]. RHINE [Car96]. RIACA [Hec94]. robust [Ke00]. roots [KFC^{+}01]. Rostock [Glo98].

San [ACM03b]. Sanitt [SRRM^{+}97]. Santander [Gut04]. SCG [ACM01]. SCG’03 [ACM03b]. Science [FGHK95, uFSGO95, SRRM^{+}97]. Scotland [Tra00]. Second [KH97, CCT04]. Semiconductors [SRRM^{+}97]. Semiconductors/High [SRRM^{+}97]. Seventeenth [ACM01]. Simon [ACM99]. Singular [GPS97]. Singularity [GPS97]. Software [Bee17, Sor97b, Sor97a, Sor97d]. Solitonen [Ano93b]. Solitons [Ano93b]. Solutions [Tec98, CCT04]. solvelib [Ano01-29]. solver [Ano01-29]. Solvers [Ber96, PZ96b, PZ96a]. Solving [BK04, RLRML04, SW99]. Some [Thi08]. Spaces [ZP06]. Spain [Gut04]. sparse [Fat03]. speed [Fat03]. Speicherverwaltung [Nau97]. Spiel [Sor97c]. St [Tra00]. stability [Ke00]. standard [Ano01-30]. Start [SRRM^{+}97]. started [Maj05a]. starting [Pop00]. state [DZ95]. statistics [Ano01-31]. stats [Ano01-31]. Steigerung [Sor96]. Sternstunden [SRRM^{+}97]. storage [Nau97]. string [Ano01-32]. stringlib [Ano01-32]. student [Ano01-33]. study [Thi00a, Thi00b]. suite [Zim95]. Summary [WRL99]. Switzerland [Lak96]. Symbolic [ACM99, ACM03a, Ber96, BCK99, Glo98, Gut04, KH97, Ke00, Lak96, Tra00, Ces97, DK98, Hec94, Nau97, SW99, Ano93b, AH09, Kra96]. symbolische [Nau97, Ano93b, Kra96]. SymmPAD [Ano93b, Ano93b]. Symposium [ACM99, ACM01, ACM03a, ACM03b, Glo98, Gut04, KH97, Lak96, Tra00]. System [Ano93b, Got94, GPS97, uRH95, uRH96, MK94, Wiw93, DK98]. Systeme [Ano93b, Nau96]. Systemen [Ano93b, Ben98, Sor96]. Systems [Ano93b, RLRML04, AH09, BK04, Ben98, GKW03, LW99, LW00, Nau96, Sor96].

Tech [SRRM^{+}97]. Technical [Vol97]. Technik [SRRM^{+}97]. Technik/Sanitt [SRRM^{+}97]. test [Zim95]. Their [RLRML04, Ben98]. Theories [SRRM^{+}97]. Theory [GPS97, Ano01-38, Ano01t]. Tool [Ano96, Wiw92, F^{+}93, Fuc94]. toolbox [Ano01-27, AH09]. tools [Ano01k, Ano01-33]. toric [Hub00]. Torun [Ano05]. transform [Ano01-34]. transforms [Ano01-34]. Transport [SRRM^{+}97]. Treatment [Ano93b]. tutorial [CO04, Fuc94, Oev00]. Tutorium [Oev00]. Two [ZP06]. Two-Level [ZP06]. type [Ano01-40, Ano01-40]. types [Ano01a, Ano01d, Ano01-29].

Uni [Ano95c]. Uni-Programm [Ano95c]. University [ACM99, Ano05, Glo98, Gut04, Tra00, Ano95c]. Up/Gambini [SRRM^{+}97]. USA [ACM03b, ACM01]. Use [Ano93b, Ben98]. User [F^{+}93, The96, Ano96, Ano01-39, Sor99]. Users [Ano05]. Using [Bee17, RLRML04, BG97, SW99]. utilities [Ano01-29].

Vancouver [ACM99]. Version [The96, Ano96, Fuc94, Oev00]. verteilte [Nau97]. Verwaltung [Sor96]. Volume [ZP06].

Web [WKZZ03]. Wester [Zim95]. Weyl [Tsa00]. Wissenschaft [uFSGO95]. within [AH09]. without [Sor97c]. WME [WKZZ03]. workshop [Car96, Hec94]. World [M^{+}95].

zur [Sor96]. Zürich [Lak96].

References

ACM:1999:IJS

[ACM99] ACM, editor. ISSAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: Proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95.

ACM:2001:PSA

[ACM01] ACM, editor. Proceedings of the Seventeenth Annual Symposium on Computational Geometry (SCG ’01): June 3–5, 2001, Medford, Massachusetts, USA. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-357-X. LCCN QA448.D38 S96 2001.

ACM:2003:IPI

[ACM03a] ACM, editor. ISSAC 2003: Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation: August 3–6, 2003, Philadelphia, PA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-641-2. LCCN QA76.95.

ACM:2003:PNA

[ACM03b] ACM, editor. Proceedings of the nineteenth annual Symposium on Computational Geometry (SCG’03): San Diego, California, USA, June 8–10, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-663-3 (paperback). LCCN QA447 .S95 2003.

Augusta:2009:PLN

[AH09] P. Augusta and Z. Hurak. Polmat library now within Symbolic Math Toolbox for Matlab in multidimensional systems computations. In IEEE, editor, International Workshop on Multidimensional (nD) Systems, 2009. nDS 2009, Aristotle University of Thessaloniki, Thessaloniki, Greece, June 29 2009–July 1 2009, pages 1–3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-2797-5. LCCN ???? IEEE catalog number CFP09612.

Anonymous:1993:PCA

[Ano93a] Anonymous. Perspectives in computer algebra. Mathematics Quarterly, 2/1993:3–6, 1993.

Anonymous:1993:SBD

[Ano93b] Anonymous. Symbolische Behandlung von Differentialgleichungen (p. 61–72), Anwendungen der Computeralgebra: Solitonen (p. 111– 116), Computeralgebra-Anwendungen bei nichtlinearen Systemen (p. 123–126), Systeme für parallele Architekturen: MuPAD (p. 180– 185), Experimentelle Systeme und Programmpakete: SymmPAD (p. 270–274). (German) [Symbolic treatment of differential equations; use of computer algebra: Solitons; computer algebra: Use in nonlinear systems; system for parallel architectures: MuPAD; experimental system and program complexes (??): SymmPAD]. In Computeralgebra in Deutschland — Bestandsaufnahme, Möglichkeiten, Perspektiven. (German) [Computer Algebra in Germany: Stocktaking, Possibilities, Perspectives]. Fachgruppe Computeralgebra der GI, DMV, GAMM, Passau und Heidelberg, Germany, 1993.

Anonymous:1995:Mb

[Ano95a] Anonymous. MuPAD. Mitteilungen des Hochschulverbandes, 1/95: 42, 1995.

Anonymous:1995:Ma

[Ano95b] Anonymous. MuPAD 1.2.2. Linux Magazin, 9/95:9, 1995.

Anonymous:1995:UPF

[Ano95c] Anonymous. Uni-Programm für Matheprobleme. (german) [University program for math problems]. Computer Zeitung, 48, November

Anonymous:1996:MUM

[Ano96] Anonymous. MuPad user’s manual: Multi-Processing Algebra Data Tool: MuPad version 1.2.2. Wiley-Teubner, Chichester, UK, 1996. ISBN 0-471-96716-5 (Wiley), 3-519-02114-5 (Teubner). xiv + 572 pp. LCCN QA76.95 .M86 1996.

Anonymous:2001:ALA

[Ano01a] Anonymous. adt — library for abstract data types. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

15 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#adt.pdf.

Anonymous:2001:ACD

[Ano01b] Anonymous. Axioms, Categories and Domains. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

69 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#domainref.pdf.

Anonymous:2001:CLC

[Ano01c] Anonymous. combinat — library for combinatorics. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25,

2001. 15 pp. URL https://www.math.utah.edu/pub/mupad/doc/
pdf_help.zip#combinat.pdf.

Anonymous:2001:DBM

[Ano01d] Anonymous. datatypes — Basic MuPAD data types. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25,

2001. 54 pp. URL https://www.math.utah.edu/pub/mupad/doc/
pdf_help.zip#datatypes.pdf.

Anonymous:2001:DLD

[Ano01e] Anonymous. detools — library for differential equations. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 28 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#detools.pdf.

Anonymous:2001:FLF

[Ano01f] Anonymous. fp — library for functional programming. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 13 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#fp.pdf.

Anonymous:2001:MMb

[Ano01g] Anonymous. From MuPAD 1.4 to MuPAD 2.0. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

59 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#changes.pdf.

Anonymous:2001:GGI

[Ano01h] Anonymous. generate — generate input to other programs. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 34 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#generate.pdf.

Anonymous:2001:GLG

[Ano01i] Anonymous. groebner — library for Gröbner bases. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 9 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#groebner.pdf.

Anonymous:2001:IID

[Ano01j] Anonymous. import — import data. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 8

pp.  URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#import.pdf.

Anonymous:2001:ITI

[Ano01k] Anonymous. intlib — tools for integration. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

10 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#intlib.pdf.

Anonymous:2001:LLLc

[Ano01l] Anonymous. linalg — library for linear algebra. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

169 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#linalg.pdf.

Anonymous:2001:LLLb

[Ano01m] Anonymous. linopt — library for linear optimization. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 47 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#linopt.pdf.

Anonymous:2001:LLLa

[Ano01n] Anonymous. listlib — library for list manipulation. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25,

2001. 15 pp. URL https://www.math.utah.edu/pub/mupad/doc/
pdf_help.zip#listlib.pdf.

Anonymous:2001:MLP

[Ano01o] Anonymous. matchlib — library for pattern matching. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 4 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#matchlib.pdf.

Anonymous:2001:MMa

[Ano01p] Anonymous. misc — miscellanea. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 9

pp.  URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#misc.pdf.

Anonymous:2001:MLD

[Ano01q] Anonymous. module — library for dynamic modules. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25,

2001. 16 pp. URL https://www.math.utah.edu/pub/mupad/doc/
pdf_help.zip#module.pdf.

Anonymous:2001:MQR

[Ano01r] Anonymous. MuPAD 2.0 Quick Reference. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

16 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#quickref.pdf.

Anonymous:2001:NLNb

[Ano01s] Anonymous. numeric — library for numerical algorithms. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 160 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#numeric.pdf.

Anonymous:2001:NLNa

[Ano01t] Anonymous. numlib — library for number theory. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

50 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#numlib.pdf.

Anonymous:2001:O

[Ano01u] Anonymous. ode. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 11 pp. URL https://www.

math.utah.edu/pub/mupad/doc/pdf_help.zip#ode.pdf.

Anonymous:2001:OLO

[Ano01v] Anonymous. orthpoly — library for orthogonal polynomials. Mu-PAD Research Group, Universität Paderborn, Paderborn, Germany,

July 25, 2001. 23 pp. URL https://www.math.utah.edu/pub/
mupad/doc/pdf_help.zip#orthpoly.pdf.

Anonymous:2001:OLF

[Ano01w] Anonymous. output — library for formatted output. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 8 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#output.pdf.

Anonymous:2001:PGL

[Ano01x] Anonymous. plot — graphics library. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 177

pp.  URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#plot.pdf.

Anonymous:2001:P

[Ano01y] Anonymous. plotlib. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 23 pp. URL https:

//www.math.utah.edu/pub/mupad/doc/pdf_help.zip#plotlib.
pdf.

Anonymous:2001:PLPb

[Ano01z] Anonymous. polylib — library for polynomial manipulation. Mu-PAD Research Group, Universität Paderborn, Paderborn, Germany,

July 25, 2001. 33 pp. URL https://www.math.utah.edu/pub/
mupad/doc/pdf_help.zip#polylib.pdf.

Anonymous:2001:PPT

[Ano01-27] Anonymous. prog — programmer’s toolbox. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

39 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#prog.pdf.

Anonymous:2001:PLPa

[Ano01-28] Anonymous. property — library for properties. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

12 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#property.pdf.

Anonymous:2001:SDT

[Ano01-29] Anonymous. solvelib — data types and utilities for the solver. MuPAD Research Group, Universität Paderborn, Paderborn, Ger-

many, July 25, 2001. 12 pp. URL https://www.math.utah.edu/
pub/mupad/doc/pdf_help.zip#solvelib.pdf.

Anonymous:2001:SL

[Ano01-30] Anonymous. The standard library. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 931

pp.  URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#stdlib.pdf.

Anonymous:2001:SLSb

[Ano01-31] Anonymous. stats — library for statistics. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

70 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#stats.pdf.

Anonymous:2001:SLSa

[Ano01-32] Anonymous. stringlib — library for string manipulation. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 11 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#stringlib.pdf.

Anonymous:2001:SET

[Ano01-33] Anonymous. student — educational tools. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

23 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#student.pdf.

Anonymous:2001:TLI

[Ano01-34] Anonymous. transform — library for integral transforms. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July

25, 2001. 8 pp. URL https://www.math.utah.edu/pub/mupad/
doc/pdf_help.zip#transform.pdf.

Anonymous:2001:APA

[Ano01-35] Anonymous. Ax — predefined axioms. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 9

pp.  URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#Ax.pdf.

Anonymous:2001:CPC

[Ano01-36] Anonymous. Cat — predefined categories. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

62 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#Cat.pdf.

Anonymous:2001:DPD

[Ano01-37] Anonymous. Dom — predefined domains. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

230 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#Dom.pdf.

Anonymous:2001:NLG

[Ano01-38] Anonymous. Network — library for graph theory. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

47 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#Network.pdf.

Anonymous:2001:PUP

[Ano01-39] Anonymous. Pref — user preferences. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001.

37 pp. URL https://www.math.utah.edu/pub/mupad/doc/pdf_
help.zip#Pref.pdf.

Anonymous:2001:TLT

[Ano01-40] Anonymous. Type — library for type checking and mathematical properties. MuPAD Research Group, Universität Paderborn, Paderborn, Germany, July 25, 2001. 79 pp. URL https://www.math.

utah.edu/pub/mupad/doc/pdf_help.zip#Type.pdf.

Anonymous:2005:MFM

[Ano05] Anonymous, editor. MathPAD 2005: The Fourth Meeting of Mu-PAD Users, 4–8 July 2005, Nicholas Copernicus University, Torun, Poland. ????, ????, 2005. ISBN ???? LCCN ???? URL http://

www.mat.uni.torun.pl/konferencje/mathpad_2005.html.

Bernardin:1999:SCJ

[BCK99] Laurent Bernardin, Bruce Char, and Erich Kaltofen. Symbolic computation in Java: an appraisement. In ACM [ACM99], pages 237– 244. ISBN 1-58113-073-2. LCCN QA76.95.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL

http://www.springer.com/us/book/9783319641096.

Benker:1998:ICS

[Ben98] Hans Benker. Ingenieurmathematik mit Computeralgebra-Systemen: AXIOM, DERIVE, MACSYMA, MAPLE, MATHCAD, MATHE-MATICA, MATHLAP und MuPAD in der Anwendung. (German) [Engineering mathematics with computer algebra systems: AXIOM, DERIVE, MACSYMA, MAPLE, MATHCAD, MATHEMATICA, MATHLAP and MuPAD in their use]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. ISBN 3-528-05673-8. xiii + 439 pp. LCCN ????

Bernardin:1996:RSS

[Ber96] Laurent Bernardin. A review of symbolic solvers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):9–20, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bernardin:1997:MMP

[Ber97] Laurent Bernardin. Maple on a massively parallel, distributed memory machine. In Kaltofen and Hitz [KH97], pages 217–222. ISBN 0-89791-951-3. LCCN QA76.58 .I5674 1997. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/266670/p217-
bernardin/.

Bachmann:1997:DEM

[BG97] Olaf Bachmann and Simon Gray. A demonstration of exchanging mathematical expressions using MP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3): 50, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Ballarin:2004:SPL

[BK04] Clemens Ballarin and Manuel Kauers. Solving parametric linear systems: an experiment with constraint algebraic programming. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(2):33–46, June 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Carriere:1996:PRW

[Car96] L. R. Oudin A. Carrière, editor. Proceedings of the 5th RHINE workshop on computer algebra. Institut Franco-Allemand de Recherches de Saint-Louis, 68300 Saint-Louis, France, April 1996. ISBN ???? LCCN ????

Chan:2004:NLS

[CCT04] L. Chan and E. S. Cheb-Terrab. Non-Liouvillian solutions for second order linear ODEs. In Gutierrez [Gut04], pages 80–86. ISBN 1-58113-827-X. LCCN ????

Cesari:1997:CCA

[Ces97] Giovanni Cesari. CALYPSO: a computer algebra library for parallel symbolic computation. In Kaltofen and Hitz [KH97], pages 204–216. ISBN 0-89791-951-3. LCCN QA76.58 .I5674 1997.

Creutzig:2004:MT

[CO04] Christopher Creutzig and W. (Walter) Oevel. MuPAD tutorial. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 3-540-22184-0. xiii + 412 pp. LCCN QA76.95 .C74 2004.

Diaz:1998:FSM

[DK98] Angel Díaz and Erich Kaltofen. FOXBOX: a system for manipulating symbolic objects in black box representation. In Gloor [Glo98], pages 30–37. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

Dupont:2003:NOP

[DLLP03] Laurent Dupont, Daniel Lazard, Sylvain Lazard, and Sylvain Petitjean. Near-optimal parameterization of the intersection of quadrics. In ACM [ACM03b], pages 246–255. ISBN 1-58113-663-3 (paperback). LCCN QA447 .S95 2003.

Drescher:1995:GBM

[DZ95] K. Drescher and P. Zimmermann. Gröbner bases in MuPAD: state and future. In Proceedings of the PoSSo workshop on software, Paris, pages 177–182. ????, ????, March 1995.

Fuchssteiner:1993:MMA

[F^{+}93] B. Fuchssteiner et al. MuPAD: multiprocessing algebra data tool: Benutzerhandbuch. (German) [MuPAD User Handbook]. Birkhäuser Verlag, Basel, Switzerland, 1993. ISBN 3-7643-2872-X.

Fateman:2003:CSP

[Fat03] Richard Fateman. Comparing the speed of programs for sparse polynomial multiplication. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(1):4–15, March 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fleischer:1995:CAS

[FGHK95] J. (Jochem) Fleischer, J. Grabmeier, F. W. Hehl, and W. Küchlin, editors. Computer Algebra in Science and Engineering: 28–31 August 1994, Bielefeld, Germany. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1995. ISBN 981-02-2319-6. LCCN QA155.7.E4 C6496 1995.

Fee:1997:PPCb

[FM97] G. J. Fee and M. B. Monagan. Problems, puzzles, challenges. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(4):48–49, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fuchssteiner:1994:MMP

[Fuc94] Benno Fuchssteiner, editor. MuPad multi processing algebra data tool: tutorial, MuPad version 1.2. Birkhäuser Verlag, Basel, Switzerland, 1994. ISBN 3-7643-5017-2 (Basel), 0-8176-5017-2 (Boston). vi + 197 pp. LCCN QA76.95 .M85 1994.

Gerdt:1997:MIB

[GB97] Vladimir P. Gerdt and Yuri A. Blinkov. Minimal involutive bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):44, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN

QA155.7.E4 C64954 2003. URL http://www.springer.com/sgw/
cda/frontpage/0,11855,1-102-22-1477871-0,00.html. Includes

CD-ROM.

Gloor:1998:IAU

[Glo98] Oliver Gloor, editor. ISSAC 98, August 13–15, 1998, University of Rostock, Germany: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

Gottheil:1994:DGB

[Got94] K. Gottheil. Differential-Gröbner-Basen im Computer-Algebra System MuPAD. (german) [Differential Gröbner bases in the computeralgebra system MuPAD]. In DMV-Jahrestagung 1994 (Duisburg, 18-24.9.94), Sektion 11 — Computeralgebra. ????, ????, 1994.

Greuel:1997:SCA

[GPS97] G.-M. Greuel, G. Pfister, and H. Schönemann. Singular 1.0: A computer algebra system for singularity theory, algebraic geometry and commutative algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):48–50, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Grossman:1995:IIC

[Gro95] Robert Grossman. 1st International IMACS Conference on Applications of Computer Algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):9–12, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: Proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-827-X. LCCN ????

Hare:2000:RCB

[Har00a] K. Hare. Rapid computation of Bernoulli and related numbers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):18–19, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Harris:2000:ANM

[Har00b] Jason Harris. Advanced notations in Mathematica. In Traverso [Tra00], pages 153–160. ISBN 1-58113-218-2. LCCN QA 155.7 .E4E88 2000.

Heck:1994:HIS

[Hec94] André Heck. Human interaction for symbolic computation: RIACA workshop, Amsterdam, March 10–11, 1994. ACM SIGCHI Bulletin, 26(3):19–22, July 1994. CODEN SGBUD4. ISSN 0736-6906.

Heckler:1997:PM

[HKM^{+}97] Christian Heckler, Oliver Kluge, Torsten Metzner, Andreas Sorgatz, and Paul Zimmermann. Parallelism in MuPAD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):48, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Huber:2000:CGF

[Hub00] Birkett Huber. Computing Gröbner fans of toric ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):19–20, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Iglio:1998:SCC

[IA98] Pietro Iglio and Giuseppe Attardi. ISSAC 98, August 13–15, 1998, University of Rostock, Germany: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation. In Gloor [Glo98], pages 62–69. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

Ke:2000:SAC

[Ke00] Nainn-Ping Ke. Symbolic and algebraic computation in robust stability analysis. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):21, March 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krishnan:2001:PEM

[KFC^{+}01] Shankar Krishnan, Mark Foskey, Tim Culver, John Keyser, and Dinesh Manocha. PRECISE: efficient multiprecision evaluation of algebraic roots and predicates for reliable geometric computation. In ACM [ACM01], pages 274–283. ISBN 1-58113-357-X. LCCN QA448.D38 S96 2001.

Kaltofen:1997:PPS

[KH97] Erich Kaltofen and Markus Hitz, editors. PASCO ’97. Proceedings of the Second International Symposium on Parallel Symbolic Computation, July 20–22, 1997, Maui, HI. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-951-3. LCCN QA76.58 .I5674 1997.

Kluge:1996:EPF

[Klu96] O. Kluge. Entwicklung einer Programmierumgebung für die Parallelverarbeitung in der Computer-Algebra. (German) [Development of a programming environment for parallel processing in computer algebra]. MuPAD Reports. Teubner, Stuttgart, Germany; Leipzig, Germany, December 1996. ISBN 3-519-02196-X.

Kai:2000:HCB

[KN00] Hiroshi Kai and Matu-Tarow Noda. Hybrid computation of bivariate rational interpolation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):20–21, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kraume:1996:SI

[Kra96] R. Kraume. Symbolische Integration. (german) [Symbolic integration]. Spektrum der Wissenschaft (German translation of Scientific American), 3:95–98, 1996. CODEN SPEKDI. ISSN 0170-2971.

Lakshman:1996:IPI

[Lak96] Y. N. Lakshman, editor. ISSAC ’96: proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation: July 24–26, 1996, Zürich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA155.7.E4.

LJS:1997:NP

[Lin97] Linux Journal Staff. New products. Linux Journal, 39es(??):??, July 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Lazard:2004:IQE

[LPP04] S. Lazard, L. M. Peñaranda, and S. Petitjean. Intersecting quadrics: an efficient and exact implementation. In Proceedings of the Twentieth Annual Symposium on Computational Geometry (SCG ’04): June 9–11, 2004, Brooklyn, New York, USA, pages 419–428. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-885-7. LCCN QA448.D38.

Lewis:1999:CPO

[LW99] Robert H. Lewis and Michael Wester. Comparison of polynomialoriented computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(4):5– 13, December 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lewis:2000:CPO

[LW00] Robert H. Lewis and Michael Wester. Comparison of polynomialoriented computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):24, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maurer:1995:EMH

[M^{+}95] Hermann A. Maurer et al., editors. Educational Multimedia and Hypermedia, 1995: proceedings of ED-MEDIA 95: World Conference on Educational Multimedia and Hypermedia, Graz, Austria, June 17–21, 1995. Association for the Advancement of Computing in Education, Charlottesville, VA, USA, 1995. ISBN 1-880094-15-0 (paperback). LCCN LB1028.5 .W67 1995.

Majewski:2002:MPC

[Maj02] Miroslaw Majewski. MuPAD Pro computing essentials. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43574-3 (softcover). vii + 456 pp. LCCN

QA76.95 .M35 2002. URL http://www.mupad.com/majewski/
edition1/.

Majewski:2004:MPC

[Maj04] Miroslaw Majewski. MuPAD Pro computing essentials. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 3-540-21943-9. xi + 538 pp. LCCN

QA76.95 M35 2004. URL http://www.mupad.com/majewski/
edition2/.

Majewski:2005:GSM

[Maj05a] Miroslaw Majewski. Getting started with MUPAD. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-28635-7 (softcover). 270 (est.) pp. URL http://www.

mupad.com/majewski/starters/.

Majewski:2005:MDN

[Maj05b] Miroslaw Majewski. MuPAD dla niecierpliwych. (Polish) [MuPAD for the impatient]. Edukacja z TI, ????, Poland, 2005. ISBN ????

200 (est.) pp. LCCN ???? URL http://www.mupad.com/mathpad/.

Fuchssteiner:1991:MOE

[mat91] Benno Fuchssteiner, Andreas Sorgatz, and Mirek Majewski, editors. mathPAD Online Edition: The Magazine of the MuPAD Research

Group, 1991. URL http://www.mupad.com/mathpad/. MuPAD Re-

search Group, Universität Paderborn, Paderborn, Germany.

McAndrew:1999:M

[McA99] Alasdair McAndrew. MuPad. Linux Journal, 63:??, July 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

http://noframes.linuxjournal.com/lj-issues/3101.html.

Morisse:1994:CAS

[MK94] K. Morisse and A. Kemper. The computer algebra system MuPAD. Euromath Bulletin, 1(2):95–102, 1994.

Morisse:1995:NDM

[MO95] K. Morisse and G. Oevel. New developments in MuPAD. In Fleischer et al. [FGHK95], pages 44–56. ISBN 981-02-2319-6. LCCN QA155.7.E4 C6496 1995.

Murao:2000:AAE

[Mur00] Hirokazu Murao. 2W -ary algorithm for extended problem of integer GCD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):25, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Naundorf:1995:PM

[Nau95] H. Naundorf. Parallelism in MuPAD. In Michael Wester, Stanley Steinberg, and Michael Jahn, editors, Electronic Proceedings of the 1st International IMACS Conference on Applications of Computer Algebra, http://math.unm.edu/ACA//1995Proceedings/MainPage.html. ????, ????, May 1995.

Naundorf:1996:DMF

[Nau96] H. Naundorf. Ein denotationales Modell für parallele objektbasierte Systeme. (German) [A denotational model for parallel object-based systems]. MuPAD Reports. Teubner, Stuttgart, Germany; Leipzig, Germany, December 1996. ISBN 3-519-02197-8.

Naundorf:1997:MVS

[Nau97] H. Naundorf. MAMMUT — Eine verteilte Speicherverwaltung für symbolische Manipulation. (German) [MAMMUT — A distributed storage management for symbolic manipulation]. MuPAD Reports. Teubner, Stuttgart, Germany; Leipzig, Germany, February 1997. ISBN 3-519-02198-6.

Oevel:2000:MTM

[Oev00] W. (Walter) Oevel, editor. MuPAD-Tutorium. MuPAD tutorial: a version and platform independent introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67546-9 (softcover). xii + 361 pp. LCCN QA76.95 .M85513 2000.

Oevel:1996:CAE

[OS96] G. Oevel and G. Siek. Computer algebra in education. In Maurer et al. [M^{+}95], page ?? ISBN 1-880094-15-0 (paperback). LCCN LB1028.5 .W67 1995.

Prieto:2000:MOA

[PDP00] Hélène Prieto, Stéphane Dalmas, and Yves Papegay. Mathematica as an OpenMath application. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):22–26, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Popov:2000:OSA

[Pop00] Bogdan A. Popov. Optimal starting approximation and iterative algorithm for inverse error function. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):25– 26, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Postel:1997:NIM

[Pos97] A. Sorgatz F. Postel. New issues in MuPAD 1.3. Technical report, CAN, Amsterdam, The Netherlands, May 1997.

Postel:1996:ROS

[PZ96a] F. Postel and P. Zimmermann. A review of the ODE solvers of Maple, Mathematica, Macsyma and MuPAD. In Carrière [Car96], page ???? ISBN ???? LCCN ???? updated version includes Axiom, Derive and Reduce.

PosZim:96

[PZ96b] F. Postel and P. Zimmermann. A review of the ODE solvers of Maple, Mathematica, Macsyma and MuPAD. In Carrière [Car96], page ???? ISBN ???? LCCN ???? URL http://math-www.

uni-paderborn.de/MuPAD/BIB/poszim96.html. updated version

includes Axiom, Derive and Reduce.

Roanes-Lozano:2004:GAS

[RLRML04] Eugenio Roanes-Lozano, Eugenio Roanes-Macías, and Luis M. Laita. The geometry of algebraic systems and their exact solving using Gröbner bases. Computing in Science and Engineering, 6(2):76–79, March/April 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://csdl.

computer.org/comp/mags/cs/2004/02/c2076abs.htm;   http:/
/csdl.computer.org/dl/mags/cs/2004/02/c2076.htm;   http:
//csdl.computer.org/dl/mags/cs/2004/02/c2076.pdf.

Sato:2000:PCB

[Sat00] Yosuke Sato. Parallel computation of Boolean Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):27–28, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sorgatz:1996:DMV

[Sor96] A. Sorgatz. Dynamische Module — Eine Verwaltung für Maschinencode-Objekte zur Steigerung der Effizienz und Flexibilität von Computeralgebra-Systemen. (German) [Dynamic modules — an administration for machine-code objects for the increase of the efficiency and flexibility of computer algebra systems], volume 1 of MuPAD Reports. Teubner, Stuttgart, Germany; Leipzig, Germany, October 1996. ISBN 3-519-02195-1. Das Konzept der Dynamischen Module sowie technische Grundlagen der Implementation in MuPAD.

Sorgatz:1997:DMSa

[Sor97a] A. Sorgatz. Dynamic modules — software integration in MuPAD. Poster Abstracts, ISSAC’97, Maui, Hawaii, USA, July 1997. URL

http://www.mupad.de/BIB/ONLINE/SORGATZ97b/.

Sorgatz:1997:DMC

[Sor97b] A. Sorgatz. Dynamic modules — the concept of software integration in MuPAD. Talk and handouts, IMACS’97, Maui, Hawaii, USA.,

July 1997. URL http://www.mupad.de/BIB/ONLINE/sorgatz97c.
ps.gz.

Sorgatz:1997:MDM

[Sor97c] A. Sorgatz. MuPAD — Dynamische Module, Spiel (fast) ohne Grenzen. (german) [MuPAD — dynamic modules, play (nearly) without limits]. Linux Magazin, 01/97, 1997.

Sorgatz:1997:DMSb

[Sor97d] Andreas Sorgatz. Dynamic modules: Software integration in Mu-PAD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):45–47, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Sorgatz:1999:DMU

[Sor99] Andreas Sorgatz. Dynamic modules: user’s manual and programming guide for MuPAD 1.4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65043-1 (pbk./CD-ROM). xxiv + 234 pp. LCCN QA76.95 .S632 1999.

Sittig:1997:LIP

[SRRM^{+}97] R. Sittig, H. Rechenberg, M. Ritsch-Marte, H. Schopper, S. Koch, H. J. Queisser, H. Nicolai, and M. Ziegler. Lübbig: The Inverse Problem/Bührke: Newtons Apfel — Sternstunden der Physik: Von Galilei bis Lise Meitner/Memel, Renneberg: Geschlechterverhältnisse in Medizin, Naturwissenschaft und Technik/Sanitt: Science as a Questioning Process/Haug, Jauho: Quantum Kinetics in Transport and Optics of Semiconductors/High Tech Start Up/ Gambini, Pullin: Loops, Knots, Gauge Theories and Quantum Gravity/MuPAD. (German) [Book reviews: …]. Physik Journal, 53(7– 8):716–719, July 1997. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Sorgatz:1999:THP

[SW99] Andreas Sorgatz and Stefan Wehmeier. Towards high-performance symbolic computing: using MuPAD as a problem solving environment. Mathematics and Computers in Simulation, 49(3):235–246, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special issue on high performance symbolic computing.

Macsyma:1998:MSI

[Tec98] Technical Staff at Macsyma, Inc. Macsyma solutions to the IS-SAC ’97 Challenge Problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):18–37, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

MuPAD:1996:MUM

[The96] The MuPAD Group. MuPAD User’s Manual — MuPAD Version 1.2.2. John Wiley and Sons, Inc., Chichester, UK, first edition, March 1996. ISBN 0-471-96716-5 (Wiley), 3-519-02114-5 (Teubner). includes a CD for Apple Macintosh and UNIX.

Thiery:2000:AIGa

[Thi00a] Nicolas M. Thiéry. Algebraic invariants of graphs: a computer aided study. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):29, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Thiery:2000:AIGb

[Thi00b] Nicolas M. Thiéry. Algebraic invariants of graphs; a study based on computer exploration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):9–20, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Thiery:2008:IAC

[Thi08] Nicolas M. Thiery. Implementing algebraic combinatorics: Some feedback from the development of MuPAD–Combinat (abstract only). ACM Communications in Computer Algebra, 42(1–2):20, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Traverso:2000:IPI

[Tra00] Carlo Traverso, editor. ISSAC 2000: Proceedings of the 2000 International Symposium on Symbolic and Algebraic Computation: 7–9 August 2000, University of St. Andrews, Scotland. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-218-2. LCCN QA 155.7 .E4E88 2000.

Tsarev:1996:ACE

[Tsa96] S. P. Tsarev. An algorithm for complete enumeration of all factorizations of a linear ordinary differential operator. In Lakshman [Lak96], pages 226–231. ISBN 0-89791-796-0. LCCN QA155.7.E4. ACM order number 505960.

Tsai:2000:WCI

[Tsa00] Harrison Tsai. Weyl closure of a D-ideal. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (1):29–30, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Oevel:1995:MBW

[uFSGO95] F. Postel und F. Schwarz G. Oevel. MuPAD in Bildung und Wissenschaft. (german) [MuPAD in education and science]. In K. P. Müller, editor, Beiträge zum Mathematikunterricht, Vorträge der 29. Bundestagung fü Didaktik der Mathematik, pages 356–359. Franz Becker Verlag, Hildesheim, Germany, 1995.

Hillebrand:1995:MLM

[uRH95] A. Sorgatz und R. Hillebrand. Mathematik unter Linux: MuPAD — Ein Computeralgebra System I. (german) [Mathematics in Linux: MuPAD — A computer algebra system, I]. Linux Magazin, 12/95: 11–14, 1995.

Hillebrand:1996:MLM

[uRH96] A. Sorgatz und R. Hillebrand. Mathematik unter Linux: MuPAD — Ein Computeralgebra System II. (german) [Mathematics in Linux: MuPAD — A computer algebra system, II]. Linux Magazin, 2/96, 3/96 (Nachdruck):60–67, 1996.

Volcheck:1997:TRCc

[Vol97] Emil Volcheck. Technical report column. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(4):50–53, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wiwianka:1992:MMA

[Wiw92] W. Wiwianka. MuPAD — multiprocessing algebra data tool. SUGinfo, 2/92:31–36, 1992.

Wiwianka:1993:CAM

[Wiw93] W. Wiwianka. Das Computer-Algebrasystem MuPAD. (german) [The computer-algebra system MuPAD]. Solaris Magazin, 1/93:3, 1993.

Wang:2003:WTW

[WKZZ03] Paul S. Wang, Norbert Kajler, Yi Zhou, and Xiao Zou. WME: towards a Web for mathematics education. In ACM [ACM03a], pages x + 273. ISBN 1-58113-641-2. LCCN QA76.95.

Wester:1999:IAC

[WRL99] Michael Wester and Eugenio Roanes-Lozano. IMACS Applications of Computer Algebra Conference 1999: Summary. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):53–57, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zimmermann:1995:WTS

[Zim95] P. Zimmermann. Wester’s test suite in MuPAD 1.2.2. Computer Algebra Nederland Nieuwsbrief, 14:53–64, April 1995.

Zapotinschi:2006:TLP

[ZP06] Radu Zapotinschi and Dorina Peter. A two-level programming approach to volume propagation in higher-dimensional spaces. In Viorel Negru et al., editors, SYNASC 2006: Eighth International Symposium on Symbolic and Numeric Algorithms for Scientific Computing: proceedings: 26-29 September, 2006, Timisoara, Romania, pages 41–46. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2740-X. LCCN Q183.9 .I65 2006. IEEE Computer Society Order number P2740.